OpenStudy (anonymous):

Which graph represents the solution set of 2|x+1/4|<9 ?? Is there any way I can solve this algebraically, without looking at the graphs?

OpenStudy (anonymous):

multiply both sides by 2 you'll get the absolute value in one side then you can set the value on the right side to minus the left side and another equation the right side equals to the plus left side.

OpenStudy (anonymous):

\[|x+1/4| = 9/2\] like this

OpenStudy (anonymous):

Then \[-(x+1/4) = 9/2 and x+1/4 = 9/2\]